# A Survey on Large Language Model based Autonomous Agents

### v5

@@ 3.2 Natural Science
Documentation and Data Management: Natural scientific research often involves the collection, organization, and synthesis of substantial amounts of literature, which requires a significant dedication of time and human resources. LLM-based agents have shown strong capabilities on language understanding and employing tools such as the internet and databases for text processing. These capabilities empower the agent to excel in tasks related to documentation and data management [115, 116, 75]. In [115], the agent can efficiently query and utilize internet information to complete tasks such as question answering and experiment planning. ChatMOF [116] utilizes LLMs to extract important information from text descriptions written by humans. It then formulates a plan to apply relevant tools for predicting the properties and structures of metal-organic frameworks. ChemCrow [75] utilizes chemistry-related databases to both validate the precision of compound representations and identify potentially dangerous substances. This functionality enhances the reliability and comprehensiveness of scientific inquiries by ensuring the accuracy of the data involved.
@@
Natural Science Education: LLM-based agents can communicate with humans fluently, often being utilized to develop agent-based educational tools [115, 117, 118, 119]. For example, [115] develops agent-based education systems to facilitate students learning of experimental design, methodologies, and analysis. The objective of these systems is to enhance students’ critical thinking and problem-solving skills, while also fostering a deeper comprehension of scientific principles. Math Agents [117] can assist researchers in exploring, discovering, solving and proving mathematical problems. Additionally, it can communicate with humans and aids them in understanding and using mathematics.  [118] utilize the capabilities of CodeX [119] to automatically solve and explain university-level mathematical problems, which can be used as education tools to teach students and researchers. CodeHelp [120] is an education agent for programming. It offers many useful features, such as setting course-specific keywords, monitoring student queries, and providing feedback to the system. EduChat [86] is an LLM-based agent designed specifically for the education domain. It provides personalized, equitable, and empathetic educational support to teachers, students, and parents through dialogue. FreeText [121] is an agent that utilizes LLMs to automatically assess students’ responses to open-ended questions and offer feedback.
@@
Social Science Political Science and Economy Out of One [29], Horton [106], Ziems et al. [105]
@@
Research Assistant Ziems et al [105], Bail et al. [114]
Documentation and Data Management ChemCrow [75], Boiko et al. [115]
@@
Natural Science Education ChemCrow [75], CodeHelp [120], Boiko et al. [115], MathAgent [117], Drori et al. [118]
Engineering CS & SE RestGPT [70], Self-collaboration [24], SQL-PALM [90], RAH [92], DB-GPT [41], RecMind [51], ChatEDA [123], InteRecAgent [124], PentestGPT [125], CodeHelp [120], SmolModels [126], DemoGPT [127], GPTEngineer [128]
Industrial Automation GPT4IA [129], IELLM [130], TaskMatrix.AI [71]
Robotics & Embodied AI ProAgent [131], LLM4RL [132], PET [133], REMEMBERER [134], DEPS [33], Unified Agent [135], SayCan [78], LMMWM [136], TidyBot [137], RoCo [93], SayPlan [31]
@@ 3.3 Engineering
Civil Engineering: In civil engineering, LLM-based agents can be used to design and optimize complex structures such as buildings, bridges, dams, roads, etc. [138] proposes an interactive framework where human architects and agents collaborate to construct structures in a 3D simulation environment. The interactive agent can understand natural language instructions, place blocks, detect confusion, seek clarification, and incorporate human feedback, showing the potential for human-AI collaboration in engineering design.
@@
Computer Science & Software Engineering: In the field of computer science and software engineering, LLM-based agents offer potential for automating coding, testing, debugging, and documentation generation [14, 18, 23, 24, 128, 126, 127]. ChatDev [18] proposes an end-to-end framework, where multiple agent roles communicate and collaborate through natural language conversations to complete the software development life cycle. This framework demonstrates efficient and cost-effective generation of executable software systems. ToolBench [14] can be used for tasks such as code auto-completion and code recommendation. MetaGPT [23] abstracts multiple roles, such as product managers, architects, project managers, and engineers, to supervise code generation process and enhance the quality of the final output code. This enables low-cost software development.  [24] presents a self-collaboration framework for code generation using LLMs. In this framework, multiple LLMs are assumed to be distinct "experts" for specific sub-tasks. They collaborate and interact according to specified instructions, forming a virtual team that facilitates each other’s work. Ultimately, the virtual team collaboratively addresses code generation tasks without requiring human intervention. LLIFT [139] employs LLMs to assist in conducting static analysis, specifically for identifying potential code vulnerabilities. This approach effectively manages the trade-off between accuracy and scalability. ChatEDA [123] is an agent developed for electronic design automation (EDA) to streamline the design process by integrating task planning, script generation, and execution. CodeHelp [120] is an agent designed to assist students and developers in debugging and testing their code. Its features include providing detailed explanations of error messages, suggesting potential fixes, and ensuring the accuracy of the code. PENTESTGPT [125] is a penetration testing tool based on LLMs, which can effectively identify common vulnerabilities, and interpret source code to develop exploits. DB-GPT [41] utilizes the capabilities of LLMs to systematically assess potential root causes of anomalies in databases. Through the implementation of a tree of thought approach, DB-GPT enables LLMs to backtrack to previous steps in case the current step proves unsuccessful, thus enhancing the accuracy of the diagnosis process.
@@
Industrial Automation: In the field of industrial automation, LLM-based agents can be used to achieve intelligent planning and control of production processes. [129] proposes a novel framework that integrates large language models (LLMs) with digital twin systems to accommodate flexible production needs. The framework leverages prompt engineering techniques to create LLM agents that can adapt to specific tasks based on the information provided by digital twins. These agents can coordinate a series of atomic functionalities and skills to complete production tasks at different levels within the automation pyramid. This research demonstrates the potential of integrating LLMs into industrial automation systems, providing innovative solutions for more agile, flexible and adaptive production processes. IELLM [130] showcases a case study on LLMs’ role in the oil and gas industry, covering applications like rock physics, acoustic reflectometry, and coiled tubing control.
@@
Robotics & Embodied Artificial Intelligence: Recent works have developed more efficient reinforcement learning agents for robotics and embodied artificial intelligence [140, 141, 142, 143, 38, 16, 132, 133, 134, 135, 78]. The focus is on enhancing autonomous agents’ abilities for planning, reasoning, and collaboration in embodied environments. In specific, [140] proposes a unified agent system for embodied reasoning and task planning. In this system, the authors design high-level commands to enable improved planning while propose low-level controllers to translate commands into actions. Additionally, one can leverage dialogues to gather information [141] to accelerate the optimization process.  [142, 143] employ autonomous agents for embodied decision-making and exploration. To overcome the physical constraints, the agents can generate executable plans and accomplish long-term tasks by leveraging multiple skills. In terms of control policies, SayCan [78] focuses on investigating a wide range of manipulation and navigation skills utilizing a mobile manipulator robot. Taking inspiration from typical tasks encountered in a kitchen environment, it presents a comprehensive set of 551 skills that cover seven skill families and 17 objects. These skills encompass various actions such as picking, placing, pouring, grasping, and manipulating objects, among others. TidyBot [137] is an embodied agent designed to personalize household cleanup tasks. It can learn users’ preferences on object placement and manipulation methods through textual examples.
@@
To promote the application of LLM-based autonomous agents, researchers have also introduced many open-source libraries, based on which the developers can quickly implement and evaluate agents according to their customized requirements [81, 144, 145, 146, 19, 147, 148, 149, 150, 151, 152, 153, 154, 127, 155, 156, 157]. For example, LangChain [149] is an open-source framework that automates coding, testing, debugging, and documentation generation tasks. By integrating language models with data sources and facilitating interaction with the environment, LangChain enables efficient and cost-effective software development through natural language communication and collaboration among multiple agent roles. Based on LangChain, XLang [147] comes with a comprehensive set of tools, a complete user interface, and support three different agent scenarios, namely data processing, plugin usage, and web agent. AutoGPT [81] is an agent that is fully automated. It sets one or more goals, breaks them down into corresponding tasks, and cycles through the tasks until the goal is achieved. WorkGPT [150] is an agent framework similar to AutoGPT and LangChain. By providing it with an instruction and a set of APIs, it engages in back-and-forth conversations with AI until the instruction is completed. GPT-Engineer [128], SmolModels [126] and DemoGPT [127] are open-source projects that focus on automating code generation through prompts to complete development tasks. AGiXT [146] is a dynamic AI automation platform designed to orchestrate efficient AI command management and task execution across many providers. AgentVerse [19] is a versatile framework that facilitates researchers in creating customized LLM-based agent simulations efficiently. GPT Researcher [152] is an experimental application that leverages large language models to efficiently develop research questions, trigger web crawls to gather information, summarize sources, and aggregate summaries. BMTools [153] is an open-source repository that extends LLMs with tools and provides a platform for community-driven tool building and sharing. It supports various types of tools, enables simultaneous task execution using multiple tools, and offers a simple interface for loading plugins via URLs, fostering easy development and contribution to the BMTools ecosystem.
@@ 4.1 Subjective Evaluation
Human Annotation: This evaluation method involves human evaluators directly scoring or ranking the outputs generated by various agents [105, 29, 22]. For example, in [20], the authors employ many annotators, and ask them to provide feedback on five key questions that directly associated with the agent capability. Similarly,  [159] assess model effectiveness by having human participants rate the models on harmlessness, honesty, helpfulness, engagement, and unbiasedness, subsequently comparing these scores across different models. In [79], annotators are asked to determine whether the specifically designed models can significantly enhance the development of rules within online communities.
Turing Test: This evaluation strategy necessitates that human evaluators differentiate between outputs produced by agents and those created by humans. If, in a given task, the evaluators cannot separate the agent and human results, it demonstrates that the agent can achieve human-like performance on this task. For instance, researchers in [29] conduct experiments on free-form Partisan text, and the human evaluators are asked to guess whether the responses are from human or LLM-based agent. In [20], the human evaluators are required to identify whether the behaviors are generated from the agents or real-humans. In EmotionBench [160], human annotations are collected to compare the emotional states expressed by LLM software and human participants across various scenarios. This comparison serves as a benchmark for evaluating the emotional intelligence of the LLM software, illustrating a nuanced approach to understanding agent capabilities in mimicking human-like performance and emotional expression.
@@
Table 3: For subjective evaluation, we use ① and ② to represent human annotation and the Turing test, respectively. For objective evaluation, we use ①, ②, ③, and ④ to represent environment simulation, social evaluation, multi-task evaluation, and software testing, respectively. $\checkmark$” indicates that the evaluations are based on benchmarks.
@@
Tachikuma [170] $\checkmark$ 07/2023
WebArena [171] - $\checkmark$ 07/2023
RocoBench [93] - ① ② ③ - 07/2023
@@ 4.2 Objective Evaluation
Metrics: In order to objectively evaluate the effectiveness of the agents, designing proper metrics is significant, which may influence the evaluation accuracy and comprehensiveness. Ideal evaluation metrics should precisely reflect the quality of the agents, and align with the human feelings when using them in real-world scenarios. In existing work, we can conclude the following representative evaluation metrics. (1) Task success metrics: These metrics measure how well an agent can complete tasks and achieve goals. Common metrics include success rate [22, 59, 12, 57], reward/score [22, 59, 138], coverage [16], and accuracy [18, 102, 40]. Higher values indicate greater task completion ability. (2) Human similarity metrics: These metrics quantify the degree to which the agent behaviors closely resembles that of humans. Typical examples include trajectory/location accuracy [164, 38], dialogue similarities [79, 102], and mimicry of human responses [102, 29]. Higher similarity suggests better human simulation performance. (3) Efficiency metrics: In contrast to the aforementioned metrics used to evaluate the agent effectiveness, these metrics aim to assess the efficiency of agent. Commonly considered metrics encompass the length of planning [57], the cost associated with development [18], the speed of inference [16, 38], and number of clarification dialogues [138].
@@
Protocols: In addition to the evaluation metrics, another important aspect for objective evaluation is how to leverage these metrics. In the previous work, we can identify the following commonly used evaluation protocols: (1) Real-world simulation: In this method, the agents are evaluated within immersive environments like games and interactive simulators. The agents are required to perform tasks autonomously, and then metrics like task success rate and human similarity are leveraged to evaluate the capability of the agents based on their trajectories and completed objectives [164, 22, 16, 59, 38, 138, 33, 170, 85, 166]. This method is expected to evaluate the agents’ practical capabilities in real-world scenarios. (2) Social evaluation: This method utilizes metrics to assess social intelligence based on the agent interactions in simulated societies. Various approaches have been adopted, such as collaborative tasks to evaluate teamwork skills, debates to analyze argumentative reasoning, and human studies to measure social aptitude [79, 102, 165, 34, 173]. These approaches analyze qualities such as coherence, theory of mind, and social IQ to assess agents’ capabilities in areas including cooperation, communication, empathy, and mimicking human social behavior. By subjecting agents to complex interactive settings, social evaluation provides valuable insights into agents’ higher-level social cognition. (3) Multi-task evaluation: In this method, people use a set of diverse tasks from different domains to evaluate the agent, which can effectively measure the agent generalization capability in open-domain environments [29, 165, 153, 172, 173, 85, 166]. (4) Software testing: In this method, researchers evaluate the agents by letting them conduct tasks such as software testing tasks, such as generating test cases, reproducing bugs, debugging code, and interacting with developers and external tools [163, 162, 169, 173]. Then, one can use metrics like test coverage and bug detection rate to measure the effectiveness of LLM-based agents.
@@
Benchmarks: Given the metrics and protocols, a crucial remaining aspect is the selection of an appropriate benchmark for conducting the evaluation. In the past, people have used various benchmarks in their experiments. For example, many researchers use simulation environments like ALFWorld [59], IGLU [138], and Minecraft  [16, 38, 33] as benchmarks to evaluate the agent capabilities. Tachikuma [170] is a benchmark that leverages TRPG game logs to evaluate LLMs’ ability to understand and infer complex interactions with multiple characters and novel objects. AgentBench [172] provides a comprehensive framework for evaluating LLMs as autonomous agents across diverse environments. It represents the first systematic assessment of LLMs as agents on real-world challenges across diverse domains. SocKET [165] is a comprehensive benchmark for evaluating the social capabilities of LLMs across 58 tasks covering five categories of social information such as humor and sarcasm, emotions and feelings, credibility, etc. AgentSims [34] is a versatile framework for evaluating LLM-based agents, where one can flexibly design the agent planning, memory and action strategies, and measure the effectiveness of different agent modules in interactive environments. ToolBench [153] is an open-source project that aims to support the development of powerful LLMs with general tool-use capability. It provides an open platform for training, serving, and evaluating LLMs based on tool learning. WebShop [85] develops a benchmark for evaluating LLM-based agents in terms of their capabilities for product search and retrieval. The benchmark is constructed using a collection of 1.18 million real-world items. Mobile-Env [166] is an extendable interactive platform which can be used to evaluate the multi-step interaction capabilities of LLM-based agents. WebArena [171] offers a comprehensive website environment that spans multiple domains. Its purpose is to evaluate agents in an end-to-end fashion and determine the accuracy of their completed tasks. GentBench [174] is a benchmark designed to evaluate the agent capabilities, including their reasoning, safety, and efficiency, when utilizing tools to complete complex tasks. RocoBench [93] is a benchmark with six tasks evaluating multi-agent collaboration across diverse scenarios, emphasizing communication and coordination strategies to assess adaptability and generalization in cooperative robotics. EmotionBench [160] evaluates the emotion appraisal ability of LLMs, i.e., how their feelings change when presented with specific situations. It collects over 400 situations that elicit eight negative emotions and measures the emotional states of LLMs and human subjects using self-report scales. PEB [125] is a benchmark tailored for assessing LLM-based agents in penetration testing scenarios, comprising 13 diverse targets from leading platforms. It offers a structured evaluation across varying difficulty levels, reflecting real-world challenges for agents. ClemBench [167] contains five Dialogue Games to assess LLMs’ ability as a player. E2E [175] is an end-to-end benchmark for testing the accuracy and usefulness of chatbots.
@@
In the above sections, we introduce both subjective and objective strategies for LLM-based autonomous agents evaluation. The evaluation of the agents play significant roles in this domain. However, both subjective and objective evaluation have their own strengths and weakness. Maybe, in practice, they should be combined to comprehensively evaluate the agents. We summarize the correspondence between the previous work and these evaluation strategies in Table 3.

### v6

@@ 3.2 Natural Science
Documentation and Data Management: Natural scientific research often involves the collection, organization, and synthesis of substantial amounts of literature, which requires a significant dedication of time and human resources. LLM-based agents have shown strong capabilities on language understanding and employing tools such as the internet and databases for text processing. These capabilities empower the agent to excel in tasks related to documentation and data management. In [114], the agent can efficiently query and utilize internet information to complete tasks such as question answering and experiment planning. ChatMOF [115] utilizes LLMs to extract important information from text descriptions written by humans. It then formulates a plan to apply relevant tools for predicting the properties and structures of metal-organic frameworks. ChemCrow [76] utilizes chemistry-related databases to both validate the precision of compound representations and identify potentially dangerous substances. This functionality enhances the reliability and comprehensiveness of scientific inquiries by ensuring the accuracy of the data involved.
@@
Natural Science Education: LLM-based agents can communicate with humans fluently, often being utilized to develop agent-based educational tools. For example, [114] develops agent-based education systems to facilitate students learning of experimental design, methodologies, and analysis. The objective of these systems is to enhance students’ critical thinking and problem-solving skills, while also fostering a deeper comprehension of scientific principles. Math Agents [116] can assist researchers in exploring, discovering, solving and proving mathematical problems. Additionally, it can communicate with humans and aids them in understanding and using mathematics.  [117] utilize the capabilities of CodeX [118] to automatically solve and explain university-level mathematical problems, which can be used as education tools to teach students and researchers. CodeHelp [119] is an education agent for programming. It offers many useful features, such as setting course-specific keywords, monitoring student queries, and providing feedback to the system. EduChat [87] is an LLM-based agent designed specifically for the education domain. It provides personalized, equitable, and empathetic educational support to teachers, students, and parents through dialogue. FreeText [120] is an agent that utilizes LLMs to automatically assess students’ responses to open-ended questions and offer feedback.
@@
Social Science Political Science and Economy Argyle et al. [29], Horton [105], Ziems et al. [104]
@@
Research Assistant Ziems et al. [104], Bail et al. [113]
Documentation and Data Management ChemCrow [76], ChatMOF [115], Boiko et al. [114]
@@
Natural Science Education ChemCrow [76], CodeHelp [119], Boiko et al. [114], MathAgent [116], Drori et al. [117], EduChat [87], FreeText [120]
Engineering CS & SE RestGPT [71], Self-collaboration [24], SQL-PALM [89], RAH [91], D-Bot [122], RecMind [53], ChatEDA [123], InteRecAgent [124], PentestGPT [125], CodeHelp [119], SmolModels [126], DemoGPT [127], GPTEngineer [128]
Industrial Automation GPT4IA [129], IELLM [130]
Robotics & Embodied AI ProAgent [131], LLM4RL [132], PET [133], REMEMBERER [134], DEPS [33], Unified Agent [135], SayCan [79], TidyBot [136], RoCo [92], SayPlan [31], TaPA [137], Dasgupta et al. [138], DECKARD [139], Dialogue shaping [140]
@@ 3.3 Engineering
Computer Science & Software Engineering: In the field of computer science and software engineering, LLM-based agents offer potential for automating coding, testing, debugging, and documentation generation [18, 23, 24, 128, 126, 127]. ChatDev [18] proposes an end-to-end framework, where multiple agent roles communicate and collaborate through natural language conversations to complete the software development life cycle. This framework demonstrates efficient and cost-effective generation of executable software systems. MetaGPT [23] abstracts multiple roles, such as product managers, architects, project managers, and engineers, to supervise code generation process and enhance the quality of the final output code. This enables low-cost software development.  [24] presents a self-collaboration framework for code generation using LLMs. In this framework, multiple LLMs are assumed to be distinct "experts" for specific sub-tasks. They collaborate and interact according to specified instructions, forming a virtual team that facilitates each other’s work. Ultimately, the virtual team collaboratively addresses code generation tasks without requiring human intervention. LLIFT [141] employs LLMs to assist in conducting static analysis, specifically for identifying potential code vulnerabilities. This approach effectively manages the trade-off between accuracy and scalability. ChatEDA [123] is an agent developed for electronic design automation (EDA) to streamline the design process by integrating task planning, script generation, and execution. CodeHelp [119] is an agent designed to assist students and developers in debugging and testing their code. Its features include providing detailed explanations of error messages, suggesting potential fixes, and ensuring the accuracy of the code. Pentest [125] is a penetration testing tool based on LLMs, which can effectively identify common vulnerabilities, and interpret source code to develop exploits. D-Bot [122] utilizes the capabilities of LLMs to systematically assess potential root causes of anomalies in databases. Through the implementation of a tree of thought approach, D-Bot enables LLMs to backtrack to previous steps in case the current step proves unsuccessful, thus enhancing the accuracy of the diagnosis process.
@@
Industrial Automation: In the field of industrial automation, LLM-based agents can be used to achieve intelligent planning and control of production processes. [129] proposes a novel framework that integrates LLMs with digital twin systems to accommodate flexible production needs. The framework leverages prompt engineering techniques to create LLM agents that can adapt to specific tasks based on the information provided by digital twins. These agents can coordinate a series of atomic functionalities and skills to complete production tasks at different levels. This research demonstrates the potential of integrating LLMs into industrial automation systems, providing innovative solutions for more agile, flexible and adaptive production processes. IELLM [130] showcases a case study on LLMs’ role in the oil and gas industry, covering applications like factory automation and PLC programming.
Robotics & Embodied Artificial Intelligence: Recent works have advanced the development of more efficient reinforcement learning agents for robotics and embodied artificial intelligence [138, 140, 139, 137, 38, 16, 132, 133, 134, 135, 79]. These efforts focus on enhancing autonomous agents’ capabilities in planning, reasoning, and collaboration within embodied environments. For instance, [138] proposes the Planner-Actor-Reporter paradigm for embodied reasoning and task planning. DECKARD [139] introduces the Planner-Actor-Reporter paradigm, which facilitates embodied reasoning and task planning by decoupling the agent’s planning, execution, and reporting processes. TaPA [137] constructs a multimodal dataset comprising multi-view RGB images of indoor scenes, human instructions, and corresponding plans to fine-tune LLMs. The fine-tuned models align visual perception with task planning, enabling them to generate more executable plans and significantly improving their performance in visually grounded tasks.
@@
To overcome the physical constraints, the agents can generate executable plans and accomplish long-term tasks by leveraging multiple skills. In terms of control policies, SayCan [79] focuses on investigating a wide range of manipulation and navigation skills utilizing a mobile manipulator robot. Taking inspiration from typical tasks encountered in a kitchen environment, it presents a comprehensive set of 551 skills that cover seven skill families and 17 objects. These skills encompass various actions such as picking, placing, grasping, and manipulating objects, among others. TidyBot [136] is an embodied agent designed to personalize household cleanup tasks. It can learn users’ preferences on object placement and manipulation methods through textual examples.
@@
To promote the application of LLM-based autonomous agents, researchers have also introduced many open-source libraries, based on which the developers can quickly implement and evaluate agents according to their customized requirements [82, 142, 143, 144, 19, 145, 146, 147, 148, 149, 150, 151, 152, 127, 153, 154, 155]. For example, LangChain [147] is an open-source framework that automates coding, testing, debugging, and documentation generation tasks. By integrating language models with data sources and facilitating interaction with the environment, LangChain enables efficient and cost-effective software development through natural language communication and collaboration among multiple agent roles. Based on LangChain, XLang [145] provides a comprehensive set of tools and a fully integrated user interface. It focuses on executable language grounding, enabling the conversion of natural language instructions into code or action sequences that interact seamlessly with various environments, including databases, web applications, and physical robots. AutoGPT [82] is an agent that is fully automated. It sets one or more goals, breaks them down into corresponding tasks, and cycles through the tasks until the goal is achieved. WorkGPT [148] is an agent framework similar to AutoGPT and LangChain. By providing it with an instruction and a set of APIs, it engages in back-and-forth conversations with AI until the instruction is completed. GPT-Engineer [128] and DemoGPT [127] are open-source projects that focus on automating code generation through prompts to complete development tasks. SmolModels [126] offers a family of compact language models suitable for various tasks. AGiXT [144] is a dynamic AI automation platform that efficiently manages instructions and executes complex tasks across various AI providers, integrating adaptive memory, smart features, and a versatile plugin system. AgentVerse [156] is a versatile framework that facilitates researchers in creating customized LLM-based agent simulations efficiently. GPT Researcher [150] is an experimental application that leverages LLMs to efficiently develop research questions, trigger web crawls to gather information, summarize sources, and aggregate summaries. BMTools [151] provides a platform for community-driven tool building and sharing. It supports various types of tools, enables simultaneous task execution using multiple tools, and offers a simple interface for loading plugins via URLs, fostering easy development and contribution to the BMTools ecosystem.
@@ 4.1 Subjective Evaluation
Human Annotation: This evaluation method involves human evaluators directly scoring or ranking the outputs generated by various agents [104, 29, 22]. For example, in [20], the authors engage numerous annotators by asking 25 questions that explore their abilities across five key areas directly related to agent capabilities. In [80], annotators are asked to determine whether the specifically designed models can significantly enhance the development of rules within online communities.
Turing Test: This evaluation strategy necessitates that human evaluators differentiate between outputs produced by agents and those created by humans. If, in a given task, the evaluators cannot separate the agent and human results, it demonstrates that the agent can achieve human-like performance on this task. For instance, researchers in [29] conduct experiments on free-form Partisan text, and the human evaluators are asked to guess whether the responses are from human or LLM-based agent.
@@
Table 3: For subjective evaluation, we use ① and ② to represent human annotation and the Turing test, respectively. For objective evaluation, we use ①, ②, ③, and ④ to represent real-world simulation, social evaluation, multi-task evaluation, and software testing, respectively. $\checkmark$” indicates that the evaluations are based on benchmarks.
@@
Tachikuma [168] ① ③ $\checkmark$ 07/2023
RocoBench [92] - ① ③ $\checkmark$ 07/2023
@@ 4.2 Objective Evaluation
Metrics: In order to objectively evaluate the effectiveness of the agents, designing proper metrics is significant, which may influence the evaluation accuracy and comprehensiveness. Ideal evaluation metrics should precisely reflect the quality of the agents, and align with the human feelings when using them in real-world scenarios. In existing work, we can conclude the following representative evaluation metrics. (1) Task success metrics: These metrics measure how well an agent can complete tasks and achieve goals. Common metrics include success rate [22, 60, 12, 58], reward/score [22, 60, 161], coverage [16], and accuracy/error rate [80, 18, 101, 40]. Depending on the scenario, accuracy may reflect aspects such as program executability [18] or task validity [101]. Higher values across these task success metrics indicate greater task completion ability. (2) Human similarity metrics: These metrics quantify the degree to which the agent behaviors closely resembles those of humans by emphasizing various aspects related to human traits, such as coherent [104], fluent [104], dialogue similarities with human [80] and human acceptance rate [101]. Higher similarity suggests better human simulation performance. (3) Efficiency metrics: In contrast to the aforementioned metrics used to evaluate the agent effectiveness, these metrics aim to assess the efficiency of agent. Commonly considered metrics encompass the cost associated with development [18] and training efficiency [16, 38].
@@
Protocols: In addition to the evaluation metrics, another important aspect for objective evaluation is how to leverage these metrics. In the previous work, we can identify the following commonly used evaluation protocols: (1) Real-world simulation: In this method, the agents are evaluated within immersive environments like games and interactive simulators. The agents are required to perform tasks autonomously, and then metrics like task success rate and human similarity are leveraged to evaluate the capability of the agents based on their trajectories and completed objectives [22, 16, 60, 38, 161, 33, 168, 86, 164, 12]. By simulating real-world scenarios, this approach aims to provide a comprehensive evaluation of the agents’ practical capabilities. (2) Social evaluation: This method utilizes metrics to assess social intelligence based on the agent interactions in simulated societies. Various approaches have been adopted, such as collaborative tasks to evaluate teamwork skills, debates to analyze argumentative reasoning, and human studies to measure social aptitude [80, 101, 163, 34]. These approaches analyze qualities such as coherence, theory of mind, and social IQ to assess agents’ capabilities in areas including cooperation, communication, empathy, and mimicking human social behavior. By subjecting agents to complex interactive settings, social evaluation provides valuable insights into agents’ higher-level social cognition. (3) Multi-task evaluation: In this method, people use a set of diverse tasks from different domains to evaluate the agent, which can effectively measure the agent generalization capability in open-domain environments [29, 163, 151, 169, 170, 86, 164, 12, 162]. (4) Software testing: In this method, researchers evaluate the agents by letting them conduct tasks such as software testing tasks, such as generating test cases, reproducing bugs, debugging code, and interacting with developers and external tools [160, 159, 167]. Then, one can use metrics like test coverage and bug detection rate to measure the effectiveness of LLM-based agents.
@@
Benchmarks: Given the metrics and protocols, a crucial aspect of evaluation is the selection of appropriate benchmarks. Over time, various benchmarks have been introduced to assess the capabilities of LLM-based agents across diverse domains and scenarios. Many studies employ environments such as ALFWorld [60], IGLU [161], and Minecraft [16, 38, 33] to evaluate agent capabilities in interactive and task-oriented simulations. Tachikuma [168] evaluates LLMs’ ability to infer and understand complex interactions involving multiple characters and novel objects through TRPG game logs. AgentBench [169] provides a comprehensive framework for evaluating LLMs as autonomous agents across diverse environments. It represents the first systematic assessment of LLMs as agents on real-world challenges across diverse domains. SocKET [163] is a comprehensive benchmark for evaluating the social capabilities of LLMs across 58 tasks covering five categories of social information such as humor and sarcasm, emotions and feelings, credibility, etc. AgentSims [34] is a versatile framework for evaluating LLM-based agents, where one can flexibly design the agent planning, memory and action strategies, and measure the effectiveness of different agent modules in interactive environments. ToolBench [151] focuses on assessing and enhancing language models’ ability to use tools, featuring 16,464 real-world RESTful APIs and diverse instructions tailored for single- and multi-tool scenarios. WebShop [86] develops a benchmark for evaluating LLM-based agents in terms of their capabilities for product search and retrieval, which is constructed using a collection of 1.18 million real-world items. Mobile-Env [164] serves as an extendable interactive platform designed to to evaluate the multi-step interaction capabilities of LLM-based agents. WebArena [173] offers a comprehensive website environment that spans multiple domains. Its purpose is to evaluate agents in an end-to-end fashion and determine the accuracy of their completed tasks. GentBench [171] is crafted to evaluate the agent capabilities, including their reasoning, safety, and efficiency, when utilizing tools to complete complex tasks. RocoBench [92] comprises six tasks that evaluate multi-agent collaboration across diverse scenarios, emphasizing communication and coordination strategies to assess adaptability and generalization in cooperative robotics. EmotionBench [172] evaluates the emotion appraisal ability of LLMs, i.e., how their feelings change when presented with specific situations. It collects over 400 situations that elicit eight negative emotions and measures the emotional states of LLMs and human subjects using self-report scales. PEB [125] is tailored for assessing LLM-based agents in penetration testing scenarios, comprising 13 diverse targets from leading platforms. It offers a structured evaluation across varying difficulty levels, reflecting real-world challenges for agents. ClemBench [165] contains five Dialogue Games to assess LLMs’ ability as a player. E2E [174] serves as an end-to-end benchmark for testing the accuracy and usefulness of chatbots.
@@
In the above sections, we introduce both subjective and objective strategies for LLM-based autonomous agent evaluation. The evaluation of the agents play significant roles in this domain. However, both subjective and objective evaluation have their own strengths and weakness. Maybe, in practice, they should be combined to comprehensively evaluate the agents. We summarize the correspondence between the previous work and these evaluation strategies in Table 3.